OpenStudy (anonymous):

maureen put $200 into a fund at 3% interest compounded annually. how much money will she have after 4 years?make a table to show how you got your answer

OpenStudy (anonymous):

compounded pay twice a years 4*2=8 3%=0.03 0.03/2=0.015 y=200(1+0.015)^8 y=200(1.015)^8 y=$225.30